# A First-order Generative Bilevel Optimization Framework for Diffusion Models

### v1

@@ Appendix A Additional related works
Bilevel optimization methods. Bilevel optimization has a long history that dates back to (Bracken & McGill, 1973). Recent efforts have focused on developing efficient gradient-based bilevel optimization methods with non-asymptotic convergence guarantees, inspired by works such as (Ghadimi & Wang, 2018; Ji et al., 2021; Hong et al., 2023; Chen et al., 2021). Since the gradient of the bilevel nested objective depends on the Hessian of the lower-level objective, existing literature proposed different Hessian inversion approximation methods including unrolling differentiation (Franceschi et al., 2017, 2018; Grazzi et al., 2020), implicit differentiation (Chen et al., 2021; Ghadimi & Wang, 2018; Hong et al., 2023; Pedregosa, 2016; Khanduri et al., 2021), conjugate gradients (Ji et al., 2021; Yang et al., 2021) and its warm-started single-loop versions (Arbel & Mairal, 2022; Li et al., 2022; Liu et al., 2023b; Xiao et al., 2023), and equilibrium backpropagation (Scellier & Bengio, 2017; Scellier, 2021); see (Zucchet & Sacramento, 2022) for a comparison. Among these methods, equilibrium backpropagation stands out as a fully first-order approach, valued for its balance of efficiency, robustness, and simplicity. Building on this principle, recent works have extended its applicability from the strongly convex setting to convex, nonconvex and constrained settings by reformulating the bilevel optimization problem as a single-level penalty problem and solving it via first-order approaches (Shen et al., 2023; Liu et al., 2022; Kwon et al., 2023, 2024; Chen et al., 2024; Lu & Mei, 2023; Jiang et al., 2024; Yao et al., 2024).
@@ Appendix B Background on bilevel optimization
The differentiability of the penalty problem relies on the differentiability of the value function $g^{*}(x)$, which is established through the extended Danskin theorem (Shen et al., 2023, Proposition 4). Specifically, the gradient of value function takes
@@ Appendix C Background on diffusion models
where $\text{where }~{}~{}~{}\mu_{t}(u_{t},u_{0})=\frac{\sqrt{\bar{q}_{t-1}}q_{t}}{1-\bar{q}_{t}}u_{0}+\frac{\sqrt{1-q_{t}}\left(1-\bar{q}_{t-1}\right)}{1-\bar{q}_{t}}u_{t}\stackrel{{\scriptstyle\textbf{(a)}}}{{=}}\frac{1}{\sqrt{1-q_{t}}}\left(u_{t}-\frac{q_{t}}{\sqrt{1-\bar{q}_{t}}}\delta_{t}\right)$ (23a)
[…]
where (a) is earned by reparameterizing (21) as $u_{t}(u_{0},\delta_{t})=\sqrt{\bar{q}_{t}}u_{0}+(1-\bar{q}_{t})\delta_{t}$ for $\delta_{t}\sim\mathcal{N}(\mu_{t},\sigma_{t}^{2}\mathbf{I}_{d})$. As $\mu_{t}$ is proportional to $\delta_{t}$, we can fit a neural network to proxy $\mu_{t}$ by optimizing the score matching loss in (2.2) in the following simplified form with explicit dependence on noise scheduler $q$
@@
which recovers DDPM in (25) when $\sigma_{t}=\sqrt{\left(1-\bar{q}_{t-1}\right)q_{t}/\left(1-\bar{q}_{t}\right)}$ in (23b) and without skipping. i.e. $t_{i}=t$, and the resulting deterministic model when $\sigma_{t}=0$ is called DDIM. In DDIM, time steps $\{t_{i}\}$ are selected using either linear ($t_{i}=\lfloor ci\rfloor$ for some $c$) or a quadratic ($t_{i}=\lfloor ci^{2}\rfloor$ for some $c$) strategy. With these designs, backward sampling steps of DDIM can be reduced from $1000$ in DDPM to $50-10$ (Song et al., 2021a).
@@ E.1 Guided Diffusion algorithm for Generative Optimization
[…]
@@
2:  Parameter: Strength parameters $\rho(t)$, $\{v_{n}\}_{n=0}^{N-1}$, number of iterations $N$, batch sizes $\{B_{n}\}$.
@@ E.2 Bilevel fine-tuning algorithm
¯$\bar{\nabla}\mathcal{L}_{\gamma}^{*}(\lambda)$ $=-\frac{1}{\lambda M_{0}}\sum_{m=1}^{M_{0}}r_{1}(\tilde{u}_{m})-\gamma\log\frac{1}{M_{0}}\sum_{m=1}^{M_{0}}\left[e^{\frac{r_{2}(\tilde{u}_{m})}{\lambda}}\right]+\gamma\log\frac{1}{M_{0}}\sum_{m=1}^{M_{0}}\left[e^{\frac{r_{1}(\tilde{u}_{m})/\gamma+r_{2}(\tilde{u}_{m})}{\lambda}}\right].$ (33)
@@ F.1 Application 1: fine-tuning diffusion model with bilevel entropy regularization learning
For the hyperparameter settings, we set the initial value of noise scheduler $q(t)$ to 1 and tune the entropy strength $\lambda$ using different methods. We use a batch size of 3 for the fine-tuning step and set optimization step 7 and repeat the optimization for 4 times. The prompts we used for generating these figures are mentioned in the figure 8 – 10. We compare our results against various baseline methods including grid search, random search, bayesian method, and weighted sum.
@@
$-EI(\lambda)=-\mathbb{E}[f(\lambda)-f(\lambda_{t}^{+})]$
@@ F.2 Application 2: bilevel noise scheduling learning
which covers (Chen, 2023) when $q_{\epsilon}=0$. Since $q(t)$ should be nondecreasing, we assign $q(t)=1-l(t)/l(t-1)$. With the use of parameterization, ZO perturbation will be added on $q_{s},q_{e},q_{\epsilon},q_{\tau}$ instead of directly on $q(t)$.
@@
Warm-start strategy. To further accelerate convergence, we avoid fully optimizing the penalty and lower-level problem with respect to $\theta$ for every $q$. Instead, we employ a warm-start strategy, initializing $\theta$ using its value from the previous epoch (Arbel & Mairal, 2022; Vicol et al., 2022; Sambharya et al., 2024). Empirically, this approach effectively reduces the inner loop for optimizing $\theta$ in the lower-level and penalty problems to $10$ and $1$, respectively. Moreover, only $3-4$ outer epochs are needed for optimizing $q$. Compared to the $100$ epochs required for single-level diffusion model training, this significantly enhances the computational efficiency of our method, as shown in Table 2. With only 2.5×2.5$\times$2.5 × the training time of a single-level diffusion model, the bilevel method achieves a $30\%$ improvement over the default model while using just $15\%$ of the time for Bayesian search.
@@
(a) Grid Search (FID= 127.04, CLIP= 31.97) (b) Bayesian Search (FID= 140.35, CLIP= 32.68) (c) Random Search (FID= 117.46, CLIP= 34.67)
@@
(d) Weighted Sum (FID= 108.95, CLIP= 35.87) (e) Bilevel (FID= 104.35, CLIP= 36.54) (f) $\lambda=0.01$ (FID= 403.66, CLIP= 17.58)
@@
(a) Grid Search (FID= 131.95, CLIP= 31.66) (b) Bayesian Search (FID= 128.15, CLIP= 34.16) (c) Random Search (FID= 127.91, CLIP= 36.34)
@@
(d) Weighted Sum (FID= 117.95, CLIP= 37.85) (e) Bilevel (FID= 105.72, CLIP= 39.05) (f) $\lambda=0.01$ (FID= 440.40, CLIP= 18.78)
@@
(a) Grid Search (FID= 144.23, CLIP= 31.37) (b) Bayesian Search (FID= 142.45, CLIP= 32.71) (c) Random Search (FID= 121.23, CLIP= 33.97)
@@
(d) Weighted Sum (FID= 118.91, CLIP= 34.53) (e) Bilevel (FID= 112.78, CLIP= 36.65) (f) $\lambda=0.01$ (FID= 450.34, CLIP= 16.26)

### v2

@@ Appendix A Additional related works
Bilevel optimization methods. Bilevel optimization has a long history that dates back to (Bracken & McGill, 1973). Recent efforts have focused on developing efficient gradient-based bilevel optimization methods with non-asymptotic convergence guarantees, inspired by works such as (Ghadimi & Wang, 2018; Ji et al., 2021; Hong et al., 2023; Chen et al., 2021). Since the gradient of the bilevel nested objective depends on the Hessian of the lower-level objective, existing literature proposed different Hessian inversion approximation methods including unrolling differentiation (Franceschi et al., 2017, 2018; Grazzi et al., 2020), implicit differentiation (Chen et al., 2021; Ghadimi & Wang, 2018; Hong et al., 2023; Pedregosa, 2016; Khanduri et al., 2021), conjugate gradients (Ji et al., 2021; Yang et al., 2021) and its warm-started single-loop versions (Arbel & Mairal, 2022; Li et al., 2022; Liu et al., 2023b; Xiao et al., 2023), and equilibrium backpropagation (Scellier & Bengio, 2017; Scellier, 2021); see (Zucchet & Sacramento, 2022) for a comparison. Among these methods, equilibrium backpropagation stands out as a fully first-order approach, valued for its balance of efficiency, robustness, and simplicity. Building on this principle, recent works have extended its applicability from the strongly convex setting to convex, nonconvex and constrained settings by reformulating the bilevel optimization problem as a single-level penalty problem and solving it via first-order approaches (Shen et al., 2025b; Liu et al., 2022; Kwon et al., 2023, 2024; Chen et al., 2024; Lu & Mei, 2023; Jiang et al., 2024; Yao et al., 2024).
@@ Appendix B Background on bilevel optimization
The differentiability of the penalty problem relies on the differentiability of the value function $g^{*}(x)$, which is established through the extended Danskin theorem (Shen et al., 2025b, Proposition 4). Specifically, the gradient of value function takes
@@ Appendix C Background on diffusion models
$\text{where }~~~\mu_{t}(u_{t},u_{0})=\frac{\sqrt{\bar{q}_{t-1}}q_{t}}{1-\bar{q}_{t}}u_{0}+\frac{\sqrt{1-q_{t}}\left(1-\bar{q}_{t-1}\right)}{1-\bar{q}_{t}}u_{t}\stackrel{{\scriptstyle\textbf{(a)}}}{{=}}\frac{1}{\sqrt{1-q_{t}}}\left(u_{t}-\frac{q_{t}}{\sqrt{1-\bar{q}_{t}}}\delta_{t}\right)$ (23a)
$\text{ and }\quad\sigma_{t}^{2}=\frac{1-\bar{q}_{t-1}}{1-\bar{q}_{t}}q_{t}$ (23b)
where (a) is earned by reparameterizing (21) as $u_{t}(u_{0},\delta_{t})=\sqrt{\bar{q}_{t}}u_{0}+(1-\bar{q}_{t})\delta_{t}$ for $\delta_{t}\sim\mathcal{N}(0,\mathbf{I}_{d})$. As $\mu_{t}$ is proportional to $\delta_{t}$, we can fit a neural network to proxy $\mu_{t}$ by optimizing the score matching loss in (2.2) in the following simplified form with explicit dependence on noise scheduler $q$
@@
which recovers DDPM in (25) when $\sigma_{t}=\sqrt{\left(1-\bar{q}_{t-1}\right)q_{t}/\left(1-\bar{q}_{t}\right)}$ in (23b) and without skipping. i.e. $t_{i}=t$, and the resulting deterministic model when $\sigma_{t}=0$ is called DDIM. In DDIM, time steps $\{t_{i}\}$ are selected using either linear ($t_{i}=\lfloor ci\rfloor$ for some $c$) or a quadratic ($t_{i}=\lfloor ci^{2}\rfloor$ for some $c$) strategy. With these designs, backward sampling steps of DDIM can be reduced from $1000$ in DDPM to 501050-10 (Song et al., 2021a).
@@ E.1 Guided Diffusion algorithm for Generative Optimization
where $g$ is a gradient vector associated with the reward function $r(\cdot)$ evaluated at the current sample $\tilde{u}_{t}$, $v$ is a given target reward value that increase along the optimization, $\bar{q}(t)=\exp(-\int_{0}^{t}q(s)ds)$, $h(t)=1-\bar{q}(t)$ are the mean and variance of $t$-th sample and $\rho(t)$ is the tunning parameter.
@@
2:Parameter: Strength parameters $\rho(t)$, $\{v_{n}\}_{n=0}^{N-1}$, number of iterations NN, batch sizes $\{B_{n}\}$.
@@ E.2 Bilevel fine-tuning algorithm
$\bar{\nabla}\mathcal{L}_{\gamma}^{*}(\lambda)$ $=-\frac{1}{\lambda M_{0}}\sum_{m=1}^{M_{0}}r_{1}(\tilde{u}_{m})-\gamma\log\frac{1}{M_{0}}\sum_{m=1}^{M_{0}}\left[e^{\frac{r_{2}(\tilde{u}_{m})}{\lambda}}\right]+\gamma\log\frac{1}{M_{0}}\sum_{m=1}^{M_{0}}\left[e^{\frac{r_{1}(\tilde{u}_{m})/\gamma+r_{2}(\tilde{u}_{m})}{\lambda}}\right].$ (33)
@@ F.1 Application 1: fine-tuning diffusion model with bilevel entropy regularization learning
For the hyperparameter settings, we set the initial value of noise scheduler $q(t)$ to 1 and tune the entropy strength $\lambda$ using different methods. We use a batch size of 3 for the fine-tuning step and set optimization step 7 and repeat the optimization for 4 times. The prompts we used for generating these figures are mentioned in the figure 9 – 11. We compare our results against various baseline methods including grid search, random search, bayesian method, and weighted sum.
@@
$EI(\lambda)=-\mathbb{E}[f(\lambda)-f(\lambda_{t}^{+})]$
@@ F.2 Application 2: bilevel noise scheduling learning
which covers (Chen, 2023) when $q_{\epsilon}=0$. Since $q(t)$ should be nondecreasing, we assign q(t)=1l(t)/l(t1)q(t)=1-l(t)/l(t-1). With the use of parameterization, ZO perturbation will be added on $q_{s},q_{e},q_{\epsilon},q_{\tau}$ instead of directly on $q(t)$.
@@
Warm-start strategy. To further accelerate convergence, we avoid fully optimizing the penalty and lower-level problem with respect to $\theta$ for every $q$. Instead, we employ a warm-start strategy, initializing $\theta$ using its value from the previous epoch (Arbel & Mairal, 2022; Vicol et al., 2022; Sambharya et al., 2024). Empirically, this approach effectively reduces the inner loop for optimizing $\theta$ in the lower-level and penalty problems to $10$ and $1$, respectively. Moreover, only 343-4 outer epochs are needed for optimizing $q$. Compared to the $100$ epochs required for single-level diffusion model training, this significantly enhances the computational efficiency of our method, as shown in Table 3. With only $2.5\times$ the training time of a single-level diffusion model, the bilevel method achieves a $30\%$ improvement over the default model while using just $15\%$ of the time for Bayesian search.
@@
(a) Grid Search
(FID= 127.04, CLIP= 31.97)
(b) Bayesian Search
(FID= 140.35, CLIP= 32.68)
(c) Random Search
(FID= 117.46, CLIP= 34.67)
@@
(d) Weighted Sum
(FID= 108.95, CLIP= 35.87)
(e) Bilevel
(FID= 104.35, CLIP= 36.54)
(f) $\lambda=0.01$
(FID= 403.66, CLIP= 17.58)
@@
(a) Grid Search
(FID= 131.95, CLIP= 31.66)
(b) Bayesian Search
(FID= 128.15, CLIP= 34.16)
(c) Random Search
(FID= 127.91, CLIP= 36.34)
@@
(d) Weighted Sum
(FID= 117.95, CLIP= 37.85)
(e) Bilevel
(FID= 105.72, CLIP= 39.05)
(f) $\lambda=0.01$
(FID= 440.40, CLIP= 18.78)
@@
(a) Grid Search
(FID= 144.23, CLIP= 31.37)
(b) Bayesian Search
(FID= 142.45, CLIP= 32.71)
(c) Random Search
(FID= 121.23, CLIP= 33.97)
@@
(d) Weighted Sum
(FID= 118.91, CLIP= 34.53)
(e) Bilevel
(FID= 112.78, CLIP= 36.65)
(f) $\lambda=0.01$
(FID= 450.34, CLIP= 16.26)